OpenStudy (anonymous):

Complete the blank (5+__)+(20-16)=7+17__

OpenStudy (anonymous):

is there a blank after the 17?

OpenStudy (anonymous):

yes which is 24 right

OpenStudy (anonymous):

(5+...)+(20-16)=7+17=24 (5+...)+4=24 (5+...)=20 ...=15

OpenStudy (anonymous):

so its not going to be 24=24

OpenStudy (anonymous):

sure,if you put 15 in the blank

OpenStudy (anonymous):

oh I see